OpenStudy (anonymous):

what is the principal value of cos^-1(-rad3/2)

OpenStudy (anonymous):

Explanation please.

OpenStudy (anonymous):

I too.. Ha ha ha..

OpenStudy (anonymous):

\[\large \cos^{-1}(-\theta) = \pi - \cos^{-1}(\theta)\]

OpenStudy (anonymous):

what are you trying to show me?

OpenStudy (anonymous):

So you will get : \[\cos^{-1}(-\frac{\sqrt{3}}{2}) = \pi - \cos^{-1}(\frac{\sqrt{3}}{2})\]

OpenStudy (anonymous):

So now solve for: \[\cos^{-1}(\frac{\sqrt{3}}{2}) \implies \cos^{-1}(\cos(\frac{\pi}{6}))\]

OpenStudy (anonymous):

oh i got it now

OpenStudy (anonymous):

Thank you :D

OpenStudy (anonymous):

And you know: \[\cos^{-1}(\cos(x)) = x\] Use this and go further..

OpenStudy (anonymous):

Do not forget to subtract it from \(\pi\)... You are welcome dear..

OpenStudy (anonymous):

If any doubt you can ask us anytime....

OpenStudy (anonymous):

Thank you for saying that! I definitely will, once again thank you for the help.

OpenStudy (anonymous):

You are welcome..